Prove (4) by breaking the proof into cases akin to the proof of Theorem 1.1.

x·y ≤ |x·y| = |x|·|y| for all x,y∈R. (4)

(for reference)

Theorem 1.1 (Triangle inequality).

|x+y|≤|x|+|y| forallx,y∈R. (3)

Proof. To prove (3), we consider each possible case so to be able to exploit the definition (1).Case 1: x ≥ 0, y ≥ 0. We then have by (1) that |x| = x, |y| = y, and |x + y| = x + y, and so (3) holds as an equality.
Case 2: x < 0, y < 0. We then have by (1) that |x| = −x, |y| = −y, and |x + y| = −x − y, and so again (3) holds as an equality.
Case3: x<0,y≥0,x+y≤0. Wethenhaveby(1)that|x|=−x,|y|=y,and|x+y|=−x−y. We see now

|x + y| = −x − y ≤ −x + y = |x| + |y|since −y ≤ 0 ≤ y and so (3) holds.

Case 4: x < 0, y ≥ 0, x + y ≥ 0. We now have |x| = −x, |y| = y, and |x + y| = x + y, and deduce|x + y| = x + y ≤ −x + y = |x| + |y|

2

since x < 0 < −x.
Case 5: The only other cases are where the roles of x and y in cases 3 and 4 are switched, and so (3) holds then as well.
Answer #1

x·y \le |x·y| = |x|·|y| for all x,y \in R. (4)

To prove (4), we consider each possible case.

Case 1: x \ge 0, y \ge 0 , then |x| = x, |y| = y => |x| . |y| = x . y

Also, x \ge 0, y \ge 0 => x . y \ge 0. Thus, |x·y| = x . y

So, |x·y| = x . y = |x| . |y| , and so (4) holds as an equality.


Case 2: x < 0, y < 0. then |x| = -x, |y| = -y => |x| . |y| = (-x) . (-y) = x . y

Also, x < 0, y < 0 => x . y \ge 0. Thus, |x·y| = x . y

So, |x·y| = x . y = |x| . |y| , and so (4) holds as an equality.


Case3: x<0, y \ge 0, then |x| = -x, |y| = y => |x| . |y| = (-x) . (y) = - x . y

Also, |x . y| = (-x) .y = -x . y

Thus,   |x . y| = |x| . |y|

Also, x<0, y \ge 0, => x . y < 0

As,  |x| > 0,  |y| > 0 => |x| . |y| > 0

=> x . y < 0 < |x| . |y|

Thus,

x . y < |x . y| = |x| . |y|

and so (4) holds.

Case 4: x \ge 0, y < 0,

then |x| = x, |y| = -y => |x| . |y| = (x) . (-y) = - x . y

Also, |x . y| = x . (-y) = - x. y

Thus,   |x . y| = |x| . |y|

Also,  x \ge 0, y < 0, => x . y < 0

As,  |x| > 0,  |y| > 0 => |x| . |y| > 0

=> x . y < 0 < |x| . |y|

Thus,

x . y < |x . y| = |x| . |y|

and so (4) holds.